# Teaching “Foundations of Mathematics” with the LEAN Theorem Prover

### v2

@@ 2.1 LEAN Interface
Several digital platforms, including online options, are available for programming with LEAN. For this study, LEAN was downloaded and installed for Visual Studio Code (VS-Code), [9, 3]. Once this setup is completed, a LEAN project could be initiated, which, when opened in VS-Code, appears as follows:
@@
Figure 1: The LEAN interface in VS-Code with the helpful infoview on the right side.
@@
LEAN’s automation helps the proof process by providing progress feedback, identifying mistakes, and even suggesting relevant theorems from Mathlib’s extensive collection [14]. Additional assistance is available when using LEAN in VS-Code. With the development of AI tools like ChatGPT, a code-suggestion AI called GitHub Copilot has been created and can be used directly in VS-Code. This package requires installation and login, which is available for free with a UZH account.
@@ 2.2 Tactics
(a) intro used in VS-Code.
@@
(a) exact used in VS-Code.
@@
(c) apply used in VS-Code.
@@
(c) constructor used in VS-Code.
@@
(c) constructor used in VS-Code.
@@
(c) cases’ used in VS-Code.
@@
(c) cases’ used in VS-Code.
@@
(c) symm used in VS-Code.
@@ 3.2 Data Gatheting
The idea of the interview is to cover as many topics from the lecture as possible. We see that there is a question about propositional logic, one about natural numbers and natural induction, one about the definition of an equivalence relation, and one on sets, specifically on indexed sets and operations on sets. Table 1 demonstrates how often a topic is covered in either the Thoma and Iannone paper [31], the interviews we conducted, the exercise sheets of the course ’Foundations of Mathematics’, and the final exam of that course. One can see that there is a high overlap between the topics chosen.
@@ 4.1 Exercise Sheets and the Exam
In this section, we present the mean and median scores from the exercise sheets and the final exam of the course ‘Foundations of Mathematics.’ We do this to compare the performance between LEAN and Non-LEAN students. Unlike the results from the interviews, this data includes the scores from all Non-LEAN students in the course.
@@ 4.3 Students’ General Performance
Although he is a Non-LEAN student, Nicola scores better than most students in the interview. We do not know much about him, as he did not attend the meetings, but we learn at the beginning of the interview that he is a first-semester student. There are always students who are really good at mathematics, and it should be clear that the difference between LEAN and Non-LEAN students is not that all LEAN students perform better than all Non-LEAN students. Nicola has a clean and structured proving style without ever having used LEAN. However, he solves the first exercise in the interview using sets, i.e., set theory, as seen in Figure 15. While he solves it correctly, it would have been much easier to do it using propositional logic, which he did not learn in the course ‘Foundations of Mathematics.’ In his example, we can see that no matter how good we are at something, learning something new is always beneficial.
@@ 5.3 Studies with Students - Teaching (with) LEAN
For this work, students were taught how to prove mathematical statements in LEAN. We chose participants from the ‘Foundations of Mathematics’ course at UZH. Of the 78 students, 12 showed interest, but only seven attended, with two dropping out for various reasons. The remaining five students, all first-semester mathematics students, continued with the sessions.
@@ Acknowledgments
We acknowledge the invaluable, ongoing support of the Lean community. Special thanks to Marius Furter for his clear explanations of the logic behind Lean tactics in [4], and to the Institute of Mathematics at the University of Zürich for enabling this study. Lastly, sincere thanks to all the students involved in this research; their dedication and collaboration were essential to making this study possible.

### v3

@@ 2.1 Lean Interface
Several digital platforms, including online options, are available for programming with Lean. For this study, Lean was downloaded and installed for Visual Studio Code (VS Code), [9, 3]. Once this setup is completed, a Lean project could be initiated, which, when opened in VS Code, appears as follows:
@@
Figure 1: The Lean interface in VS Code with the helpful infoview on the right side.
@@
Lean’s automation helps the proof process by providing progress feedback, identifying mistakes, and even suggesting relevant theorems from Mathlib’s extensive collection [14]. Additional assistance is available when using Lean in VS Code. With the development of AI tools like ChatGPT, a code-suggestion AI called GitHub Copilot has been created and can be used directly in VS Code. This package requires installation and login, which is available for free with a UZH account.
@@ 2.2 Tactics
(a) intro used in VS Code.
@@
(a) exact used in VS Code.
@@
(c) apply used in VS Code.
@@
(c) constructor used in VS Code.
@@
(c) constructor used in VS Code.
@@
(c) cases’ used in VS Code.
@@
(c) cases’ used in VS Code.
@@
(c) symm used in VS Code.
@@ 3.2 Data Gatheting
The idea of the interview is to cover as many topics from the lecture as possible. We see that there is a question about propositional logic, one about natural numbers and natural induction, one about the definition of an equivalence relation, and one on sets, specifically on indexed sets and operations on sets. Table 1 demonstrates how often a topic is covered in either the Thoma and Iannone paper [31], the interviews we conducted, the exercise sheets of the course “Foundations of Mathematics”, and the final exam of that course. One can see that there is a high overlap between the topics chosen.
@@ 4.1 Exercise Sheets and the Exam
In this section, we present the mean and median scores from the exercise sheets and the final exam of the course “Foundations of Mathematics”. We do this to compare the performance between Lean and Non-Lean students. Unlike the results from the interviews, this data includes the scores from all Non-Lean students in the course.
@@ 4.3 Students’ General Performance
Although he is a Non-Lean student, Nicola scores better than most students in the interview. We do not know much about him, as he did not attend the meetings, but we learn at the beginning of the interview that he is a first-semester student. There are always students who are really good at mathematics, and it should be clear that the difference between Lean and Non-Lean students is not that all Lean students perform better than all Non-Lean students. Nicola has a clean and structured proving style without ever having used Lean. However, he solves the first exercise in the interview using sets, i.e., set theory, as seen in Figure 15. While he solves it correctly, it would have been much easier to do it using propositional logic, which he did not learn in the course “Foundations of Mathematics”. In his example, we can see that no matter how good we are at something, learning something new is always beneficial.
@@ 5.3 Studies with Students - Teaching (with) Lean
For this work, students were taught how to prove mathematical statements in Lean. We chose participants from the “Foundations of Mathematics” course at UZH. Of the 78 students, 12 showed interest, but only seven attended, with two dropping out for various reasons. The remaining five students, all first-semester mathematics students, continued with the sessions.
@@ Acknowledgments
We acknowledge the invaluable, ongoing support of the Lean community. Special thanks to Marius Furter for his clear explanations of the logic behind Lean tactics in [4], and to the Institute of Mathematics at the University of Zurich for enabling this study. Lastly, sincere thanks to all the students involved in this research; their dedication and collaboration were essential to making this study possible.